OpenStudy (anonymous):

a committee of 6 is to be selected from 10 people of whom A and B are two. How many committees can be formed excluding A if B is included?

OpenStudy (wasiqss):

ecludin A means u left with 9 choices hence answer is 9C6

OpenStudy (anonymous):

incorrect the answer is 140 but i have no idea how they got that

OpenStudy (anonymous):

You mean the probability? that's 10-1 =9 so; 6/9 = 2/3

OpenStudy (anonymous):

no i don't mean the probability

OpenStudy (anonymous):

oOh, I'm sorry

OpenStudy (slaaibak):

2 * 1C0 * 1C1 * 8C5 + 1C0*1C0 * 8C6 = 140 First case: A is chosen and B is not. Second case: B is chosen and A is not Third case: Neither A nor B is chosen. Second case is similar to first, that's why I multiplied with 2. Ask if you don't understand

OpenStudy (anonymous):

Explain please.

OpenStudy (slaaibak):

If A and B may not be together on the committee, it leaves us with three possibilities: 1. A is on, B is not. 1 Choose 0 (Do not choose B) * 1 Choose 1 (Do choose A) * 8 Choose 5 (5 Remaining people on committee) 2. B is on, A is not. 1 Choose 0 (Do not choose A) * 1 Choose 1 (Do choose B) * 8 Choose 5 (5 Remaining people on committee) 3. B and A are NOT on the committee. 1 Choose 0 (Do not choose B) * 1 Choose 0 (Do not choose A) * 8 Choose 6 (6 Remaining people on committee)

OpenStudy (anonymous):

So it's basically: 1. 1 guy in 0 spots (Not choose B), 1 guy in 1 spot (Choose A), 8 guys in 5 spots (The remaining). And so on

OpenStudy (slaaibak):

Yeah

OpenStudy (anonymous):

2 * 1C0 * 1C1 * 8C5 + 1C0*1C0 * 8C6 = 140 i don't understand this?

OpenStudy (slaaibak):

I explained above